A projectile is fired at an angle of /4 to the horizontal with an initial speed of 100 m/s.
Find out parametric equations for the projectile motion using time as parameter. Find out the
height of the object after time t=2 sec.
1
Expert's answer
2020-11-19T09:20:30-0500

answer

horizontal velocity ="v_x=vcos\theta"

vertical velocity("v_y)=vsin\theta"

vertical displacement

"y=vsin\theta t-\frac{1}{2}gt2" ........eq.1


horizontal displacement


"x=vcos\theta t"


"t=\frac{x}{vcos\theta}" .......eq.2

by equation 1 and 2

"y=tan\theta x-\frac{gx^2}{2v^2cos^2\theta}"


this is the parametric equation of projectile.


maximum height can be given as

"h_m=\frac{v^2sin^2\theta}{2g}"

by putting the value


"h_m=\frac{(100)^2sin^24^o}{2(9.81)}=1.83m"